# Exact Expressive Power of Transformers with Padding

### v1

@@ Abstract
Chain of thought is a natural inference-time method for increasing the computational power of transformer-based large language models (LLMs), but comes at the cost of sequential decoding. Are there more efficient alternatives to expand a transformer’s expressive power without adding parameters? We consider transformers with padding tokens as a form of parallelizable test-time compute. We show that averaging-hard-attention, masked-pre-norm transformers with polynomial padding converge to precisely the class $\mathsf{TC}^{0}$ of extremely parallelizable problems. While the $\mathsf{TC}^{0}$ upper bound was known, proving a matching lower bound had been elusive. Further, our novel analysis reveals the precise expanded power of padded transformers when coupled with another form of inference-time compute, namely dynamically increasing depth via looping. Our core technical contribution is to show how padding helps bring the notions of complete problems and reductions, which have been a cornerstone of classical complexity theory, to the formal study of transformers. Armed with this new tool, we prove that padded transformers with $O(\log^{d}n)$ looping on inputs of length $n$ recognize exactly the class $\mathsf{TC}^{d}$ of moderately parallelizable problems. Thus, padding and looping together systematically expand transformers’ expressive power: with polylogarithmic looping, padded transformers converge to the class $\mathsf{NC}$, the best that could be expected without losing parallelism (unless $\mathsf{NC}=\mathsf{P}$). Our results thus motivate further exploration of padding and looping as parallelizable alternatives to chain of thought.
@@ 1 Introduction
Due to the computational limitations of transformers (Merrill and Sabharwal, 2023a; Strobl et al., 2024; Chiang, 2025), solving complex reasoning problems requires extending their computational power at inference time, typically by allowing models to generate long chains of thought (CoT) before their outputs (Wei et al., 2022; Nye et al., 2022). Theoretical work has shown how CoT expands the expressive power of transformers to sequential problems outside $\mathsf{TC}^{0}$, but it also sacrifices parallelism (Merrill and Sabharwal, 2023b; Li et al., 2024), making inference slow. Are there alternative inference-time compute approaches that increase expressive power while preserving parallelism?
@@
One method for parallelizable inference-time compute with LLMs is using padding tokens rather than CoT. Padding can be understood as a restriction of CoT where the tokens on the chain are restricted to some “blank” symbol rather than tokens generated by the LLM. Since all the input tokens are known in advance, this makes padding more parallelizable than CoT. There have been some attempts to make padding practical with mixed results (Goyal et al., 2024; Pfau et al., 2024), but it is not fully understood. Specifically, while it is known that padded transformers are in $\mathsf{TC}^{0}$, it has been open and elusive whether they can solve all problems in $\mathsf{TC}^{0}$ or even the smaller class $\mathsf{AC}^{0}$ (Pfau et al., 2024).
@@
Our first contribution in this work is exactly characterizing the expressive power of transformers with polynomial padding as $\mathsf{TC}^{0}$, answering an open question (Pfau et al., 2024). The result emerges through a finer-grained analysis of transformers in terms of string logics (Merrill and Sabharwal, 2023c; Chiang et al., 2023; Yang et al., 2024; Yang and Chiang, 2024). Thinking in terms of logic, we show that $n^{k}$ padding tokens give transformers enough “storage space” to resolve any first-order majority logic formula over $k$ variables. This suffices to capture all of $\mathsf{TC}^{0}$ (Mix Barrington et al., 1990), giving an exact characterization of the expressive power of such transformers.
This first result, however, does not clarify whether padded transformers can be minimally extended to gain expressivity beyond $\mathsf{TC}^{0}$. To address this, we consider the combination of padding with looping, i.e., repeating a block of layers dynamically as a function of input length (Dehghani et al., 2019; Giannou et al., 2023). This can be understood as a form of inference-time compute where padding controls the computation width and looping controls the computation depth—crucially, without adding any parameters. If the number of repetitions is minimal (e.g., sublinear) in sequence length, this transformer model remains highly parallelizable relative to CoT (Merrill and Sabharwal, 2024).
We extend our first result to give an exact characterization of the expressive power of padded transformers with minimal looping. Specifically, transformers with polynomial padding and $O(\log^{d}n)$ looping recognize exactly $\mathsf{TC}^{d}$. This means that log or polylog-looped transformers can solve many problems outside $\mathsf{TC}^{0}$ under standard complexity conjectures, including boolean formula evaluation, iterated matrix multiplication, graph connectivity, and context-free language recognition. It also follows that polylog-looped padded transformers converge in expressive power to $\mathsf{NC}$, the ceiling that could be expected while preserving parallelism under standard complexity conjectures.
Our first result about the expressive power of fixed-depth padded transformers provides crucial insight for our extended results about looped padded transformers (cf. Figure 1), by allowing, for the first time, the use of familiar tools of reductions and complete problems from classical complexity theory in the analysis of transformers. First, the ability to express $\mathsf{TC}^{0}$ implies that padded transformers can implement $\mathsf{FO}$ reductions. We then scaffold the ability to implement $\mathsf{FO}$ reductions to show that transformers can also implement $\mathsf{L}$ reductions, via known results about transformers’ ability to implement graph connectivity, an $\mathsf{NL}$-complete problem under $\mathsf{FO}$ reductions. We then use this to show that $O(\log^{d}n)$ looping recognizes exactly $\mathsf{L}$-uniform $\mathsf{TC}^{d}$, via the $\mathsf{TC}^{d}$ circuit evaluation problem, which is complete for this class under $\mathsf{L}$ reductions. This establishes the main result.
Overall, our results provide an exact characterization of the expressive power of padded fixed-depth transformers as $\mathsf{TC}^{0}$, answering an open question (Pfau et al., 2024). Further, we exactly characterize the expressive power of polylog-looped padded transformers, which reveals that padding and looping dramatically extend the expressive power of transformers under standard complexity conjectures. We take these results to motivate empirical investigation of padding and looping as forms of inference-time compute that are more parallelizable than standard CoT.
@@
Figure 1: Summary of our core results: exact charactizations of the expressive power of padded fixed-depth transformers and $O(\log^{d}n)$-depth looped, padded transformers. Via Theorem 1, we show that $\mathsf{AHAT}^{0}_{*}=\mathsf{FO}\textrm{-uniform}\;\mathsf{TC}^{0}$. Combining Theorem 1 with Theorems 2 and 3, we show that, for $d\geq 1$, $\mathsf{AHAT}^{d}_{*}=\mathsf{L}\textrm{-uniform}\;\mathsf{TC}^{d}$. More finegrained results are discussed in each section.
@@ 2.1 Averaging-Hard-Attention Transformers
Following previous work, we analyze a model of transformers that is slightly idealized compared to standard soft-attention transformers. Specifically, we analyze transformers with averaging hard attention (AHAT; also called “saturated” by Merrill et al., 2022) and masked pre-norm (Merrill and Sabharwal, 2023b). AHAT attention heads only attend to the values that maximize the attention score; in the case of ties, the head returns a uniform average of all values with tied scores. Masked pre-norm means that the transformer can apply a linear projection before pre-norm at the beginning of each sublayer, which is useful for reading individual values from the residual stream without interference from other values. Overall, we believe these idealizations are minimal changes that make it easier to implement algorithmic constructions that generalize to any sequence length.
More formally, we define the transformer sublayers in this AHAT model following Merrill and Sabharwal (2024). Each sublayer uses masked pre-norm (Xiong et al., 2020; Merrill and Sabharwal, 2023b), reading input $\mathbf{z}_{i}=\mathsf{layer\_norm}(\mathbf{m}\mathbf{h}_{i})$, where $\mathbf{h}_{i}$ is the previous sublayer output and layer-norm can be standard layer-norm (Ba et al., 2016) or RMS norm (Zhang and Sennrich, 2019). The sublayer then outputs $\delta_{1},\ldots,\delta_{n}$, and the residual stream is updated as $\mathbf{h}^{\prime}_{i}=\mathbf{h}_{i}+\delta_{i}$.
@@
A transformer defines a function $\Sigma^{*}\to\Sigma$ or $\Sigma^{*}\to\{0,1\}$ if we add a linear head to the final layer and take its argmax as the output. We say that a transformer $T:\Sigma\to\{0,1\}$ recognizes a language $L$ (with beginning-of-sequence token \$currency-dollar\\$\$) if, for any $w\in\Sigma^{*}$, $T(\$w)=1$ if and only if $w\in L$.
@@
In practice, transformers use low precision (relative to $\lvert w\rvert$) during training and inference. To justify that our constructions do not unrealistically rely on too much precision, we make sure that they use at most $O(c\log n)$ precision, following previous work (Merrill and Sabharwal, 2023a).
@@
Our constructions will use the layer-norm hash representation (Merrill and Sabharwal, 2023b, 2024) for query-key matching with positive integer values.
@@
Given $z\in\mathbb{N}$, its layer-norm hash is the vector $\langle z,1,-z,-1\rangle/\sqrt{z^{2}+1}$
@@ 2.2 Padded and Looped AHATs
We assume a looped transformer model as previously defined by Merrill and Sabharwal (2024):
@@
Looped transformers provide a way of dynamically scaling width at test time by looping a block of layers. In addition, we also consider padding (Pfau et al., 2024) as a simple way to increase width:
@@
Let $d,k\in\mathbb{Z}_{\geq 0}$. $\mathsf{AHAT}^{d}_{k}$ is the class of languages recognizable by a causally masked universal transformer with masked pre-norm, a beginning-of-sequence token, no position embedding, $O(\log^{d}n)$ depth, and $O(n^{k})$ padding tokens. Further, $\mathsf{AHAT}^{d}_{*}=\bigcup_{k=0}^{\infty}\mathsf{AHAT}^{d}_{k}$, $\mathsf{AHAT}^{*}_{k}=\bigcup_{d=0}^{\infty}\mathsf{AHAT}^{d}_{k}$, and $\mathsf{AHAT}^{*}_{*}=\bigcup_{d=0}^{\infty}\mathsf{AHAT}^{d}_{*}$.
$\mathsf{AHAT}^{d}_{k}$ uses causal masking with no positional embeddings. In contrast, some of our constructions will be for unmasked transformers with $i/n$ position embeddings (standard in theoretical analysis; Merrill and Sabharwal, 2023a). We will write $\mathsf{uAHAT}^{d}_{k}$ to signify these classes. Similarly, we will occasionally consider mixed-masked transformers where some attention heads use causal masking and others do not. We will denote these classes as $\mathsf{mAHAT}^{d}_{k}$.
With some abuse of notation, we will use $\mathsf{AHAT}^{d}_{k}$ to denote both the class of languages as defined above as well as the corresponding class of transformer models, i.e., transformers with $O(\log^{d}n)$ and $(n^{k})$ padding tokens. The distinction should be clear from the context.
@@ 2.3 Circuit Complexity
We define the circuit complexity classes $\mathsf{AC}^{d}$ and $\mathsf{TC}^{d}$ in the standard way (Arora and Barak, 2009; Strobl et al., 2024). A circuit family is a mapping $\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ from input lengths $n$ to circuits that take $n$ inputs. $\mathsf{AC}^{d}$ is the class of language that can be recognized by polynomial-size, fixed-depth circuit families with unbounded-arity AND/OR gates. $\mathsf{TC}^{d}$ is the same class but augmented with unbounded-arity MAJ gates that take a majority vote over their input bits.
It will often be useful to talk about uniform variants of these classes. A $\mathsf{X}$-uniform circuit family obeys that constraint that circuit $C_{n}$ for input size $n$ can be “built” from the input string $1^{n}$ using computation in $\mathsf{X}$ (For a more formal definition, see Strobl et al., 2024). We will consider two standard types of uniformity conditions: $\mathsf{FO}$ uniformity (which is equivalent to $\mathsf{DLOGTIME}$ uniformity; Mix Barrington et al., 1990) and the weaker $\mathsf{L}$ uniformity (i.e., log-space uniformity).
Finally, we will also use the notion of completeness for circuit classes. Informally, an $\mathsf{X}$-complete problem is a problem in $\mathsf{X}$ to which any other problem in $\mathsf{X}$ can be mapped via some simple reduction $\mathsf{R}$. Whether a problem is complete depends on the notion of reduction used. See Definition 9 for a more formal definition of reductions, which makes completeness fully defined.

### v2

@@ Abstract
Chain of thought is a natural inference-time method for increasing the computational power of transformer-based large language models (LLMs), but comes at the cost of sequential decoding. Are there more efficient alternatives to expand a transformer’s expressive power without adding parameters? We consider transformers with padding tokens as a form of parallelizable test-time compute. We show that averaging-hard-attention, masked-pre-norm transformers with polynomial padding recognize precisely the class $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ of extremely parallelizable problems. While the $\mathsf{TC}^{0}$ upper bound was known, proving a matching lower bound had been elusive. Further, our novel analysis reveals the precise expanded power of padded transformers when coupled with another form of inference-time compute, namely dynamically increasing depth via looping. Our core technical contribution is to show how padding helps bring the notions of complete problems and reductions, which have been a cornerstone of classical complexity theory, to the formal study of transformers. Armed with this new tool, we prove that padded transformers with $\mathrm{O}(\log^{d}n)$ looping on inputs of length $n$ recognize exactly the class $\mathsf{FO}$-uniform $\mathsf{TC}^{d}$ of moderately parallelizable problems. Thus, padding and looping together systematically expand transformers’ expressive power: with polylogarithmic looping, polynomially padded transformers recognize precisely the class $\mathsf{FO}\textrm{-uniform}\ \mathsf{NC}$, the best that could be expected without losing parallelism (unless $\mathsf{NC}=\mathsf{P}$). Our results thus motivate further exploration of padding and looping as parallelizable alternatives to chain of thought for test-time compute.
@@ 1 Introduction
Due to the computational limitations of transformers (merrill-sabharwal-2023-parallelism; strobl-etal-2024-formal; chiang2025transformers), solving complex reasoning problems requires extending their computational power at inference time, typically by allowing models to generate long chains of thought (CoT) before their outputs (wei2022chain; nye2022show). Theoretical work has shown how CoT expands the expressive power of transformers to sequential problems outside the class $\mathsf{TC}^{0}$ of highly parallelizable problems, but it also sacrifices parallelism (merrill2023cot; li2024chain), making inference slow. Are there alternative inference-time compute approaches that increase expressive power while preserving parallelism?
@@
One method for parallelizable inference-time compute with LLMs is using padding tokens rather than CoT. Padding can be understood as restricted CoT where the tokens on the chain are restricted to some “blank” symbol rather than tokens generated by the LLM. Since all the input tokens are known in advance, padding is more parallelizable than CoT. There have been some attempts to make padding practical with mixed results (goyal2024think; pfau-2024-think), but it is not fully understood. Specifically, while it is known that padded transformers remain in $\mathsf{TC}^{0}$, it has been open and elusive whether they can solve all problems in $\mathsf{TC}^{0}$ or even the smaller class $\mathsf{AC}^{0}$ (pfau-2024-think).
@@
Our first contribution is exactly characterizing the expressive power of transformers [^1] Our formal model, detailed in Section 2, assumes a fully uniform transformer with fixed parameters, fixed width, logarithmic or polynomial precision (showing their equivalence in this setting), masked pre-norm, and causal masking (Theorem 1 also applies to unmasked transformers, showing their equivalence under padding). with polynomial padding as $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$, answering an open question (pfau-2024-think). The result emerges through a finer-grained analysis of transformers in terms of string logics (merrill2023logic; chiang-2023-tighter; yang2024masked; yang2024counting). Thinking in terms of logic, we show that $n^{k}$ padding tokens give transformers enough “storage space” to resolve any first-order majority logic formula over $k$ variables. This suffices to capture all of $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ (barrington-1990-uniformity), giving an exact characterization of the expressive power of such transformers.
@@
Figure 1: Summary of core results: exact characterizations of the expressive power of $O(\log^{d}n)$-depth looped AHATs with padding, for $d\geq 0$. Theorem˜1 shows that $\mathsf{AHAT}^{0}_{*}=\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$. Theorem˜2 extends this to show that, for $d\geq 0$, $\mathsf{AHAT}^{d}_{*}=\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$. In the process of obtaining these results, we also found the novel circuit complexity result that, for any $d\geq 1$, $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}=\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$ Theorem˜3. Thus, for $d\geq 1$, $\mathsf{AHAT}^{d}_{*}=\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$.
This first result, however, does not clarify whether padded transformers can be minimally extended to gain expressivity beyond $\mathsf{TC}^{0}$. To address this, we consider the combination of padding with looping, i.e., repeating a block of layers dynamically as a function of input length, also referred to as universal transformers (dehghani2018universal; giannou-2023-looped). This can be understood as a form of inference-time compute where padding controls the computation width and looping controls the computation depth—crucially, without adding any parameters. If the number of repetitions is minimal (e.g., sublinear) in sequence length, this transformer model remains highly parallelizable relative to CoT (merrill-2025-little).
Extending this result, we show that, for $d\geq 1$, transformers with polynomial padding and $\mathrm{O}(\log^{d}n)$ looping recognize exactly $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$. Thus log or polylog-looped transformers can solve many problems outside $\mathsf{TC}^{0}$ under standard complexity conjectures, including boolean formula evaluation, iterated matrix multiplication, graph connectivity, and context-free language recognition. It also follows that polylog-looped padded transformers converge in expressive power to $\mathsf{FO}\textrm{-uniform}$ $\mathsf{NC}$, the ceiling that could be expected while preserving parallelism under standard complexity conjectures.
Our first result about the expressive power of fixed-depth padded transformers provides crucial insight for our extended results about looped padded transformers (cf. Figure˜1), by allowing, for the first time, the use of familiar tools of reductions and complete problems from classical complexity theory in the analysis of transformers. First, the ability to express $\mathsf{TC}^{0}$ implies that padded transformers can implement $\mathsf{FO}$ reductions. We then scaffold the ability to implement $\mathsf{FO}$ reductions to show that transformers can also implement $\mathsf{L}$ reductions, via known results about transformers’ ability to implement graph connectivity, an $\mathsf{NL}$-complete problem under $\mathsf{FO}$ reductions. We then use this to show that $\mathrm{O}(\log^{d}n)$ looping recognizes exactly $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$, via the “wide” $\mathsf{TC}^{d}$ circuit evaluation problem, which we prove to be complete for this class under $\mathsf{L}$ reductions.
Finally, we prove a novel circuit complexity result that $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}=\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$ for any $d\geq 1$. This was obtained in our analysis of looped transformers with $d\geq 1$. Besides being potentially of independent interest in circuit complexity, this allows us to see that, for $d\geq 1$, $\mathrm{O}(\log^{d}n)$ looping allows recognizing not just $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ but also $\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$ (because they are the same).
Overall, our results provide an exact characterization of the expressive power of padded fixed-depth transformers as $\mathsf{TC}^{0}$, answering an open question (pfau-2024-think). Further, we exactly characterize the expressive power of polylog-looped padded transformers, which reveals that padding and looping dramatically extend the expressive power of transformers under standard complexity conjectures. We take these results to motivate empirical investigation of padding and looping as forms of inference-time compute that are more parallelizable than standard CoT.
@@ 2.1 Averaging-Hard-Attention Transformers
Following previous work, we analyze a model of transformers that is slightly idealized compared to standard soft-attention transformers. Specifically, we analyze transformers with averaging hard attention (AHAT; also called “saturated” by merrill-etal-2022-saturated) and masked pre-norm (merrill2023cot). AHAT attention heads only attend to the values that maximize the attention score; in the case of ties, the head returns a uniform average of all values with tied scores. Masked pre-norm means that the transformer can apply a linear projection before pre-norm at the beginning of each sublayer, which is useful for reading individual values from the residual stream without interference from other values. Overall, we believe these idealizations are minimal changes that make it easier to implement algorithmic constructions that generalize to any sequence length.
More formally, we define the transformer sublayers in this AHAT model following merrill-2025-little. Each sublayer uses masked pre-norm (xiong-2020-layer; merrill2023cot), reading input $\mathbf{z}_{i}=\mathsf{layer\_norm}(\mathbf{M}\mathbf{h}_{i})$, where $\mathbf{h}_{i}$ is the previous sublayer output, $\mathbf{M}$ is some matrix, and layer-norm can be standard layer-norm (ba2016layernormalization) or RMS norm (zhang2019rms). The sublayer outputs $\delta_{1},\ldots,\delta_{n}$, and the residual stream is updated as $\mathbf{h}^{\prime}_{i}=\mathbf{h}_{i}+\delta_{i}$.
@@
[…]
@@
In Appendix˜A, we formalize logarithmic (merrill-sabharwal-2023-parallelism) and polynomial precision (chiang2025transformers) datatypes. All our constructions go through with either datatype, showing that logarithmic and polynomial-precision looped padded transformer classes are, in fact, identical.
@@
Our constructions will use the layer-norm hash representation (merrill2023cot; merrill-2025-little) for query-key matching with positive integer values.
@@
Given $z\in\mathbb{N}$, its layer-norm hash is the vector $\langle z,1,-z,-1\rangle/\sqrt{2z^{2}+2}$ .
@@ 2.2 Padded and Looped AHATs
We assume a looped transformer model as previously defined by merrill-2025-little:
@@
Looped transformers provide a way of dynamically scaling width at test time by looping a block of layers. In addition, we also consider padding (pfau-2024-think) as a simple way to increase width:
@@
Let $d,k\in\mathbb{Z}_{\geq 0}$. $\mathsf{AHAT}^{d}_{k}$ is the class of languages recognizable by a causally masked looped transformer with masked pre-norm, a beginning-of-sequence token, no position embedding, $\mathrm{O}(\log^{d}n)$ depth, and $\mathrm{O}(n^{k})$ padding tokens. Further, $\mathsf{AHAT}^{d}_{*}=\bigcup_{k=0}^{\infty}\mathsf{AHAT}^{d}_{k}$, $\mathsf{AHAT}^{*}_{k}=\bigcup_{d=0}^{\infty}\mathsf{AHAT}^{d}_{k}$, and $\mathsf{AHAT}^{*}_{*}=\bigcup_{d=0}^{\infty}\mathsf{AHAT}^{d}_{*}$.
[…]
With some abuse of notation, we will use $\mathsf{AHAT}^{d}_{k}$ to denote both the class of languages as defined above as well as the corresponding class of transformer models, i.e., transformers with $\mathrm{O}(\log^{d}n)$ and $(n^{k})$ padding tokens. The distinction should be clear from the context.
@@ 2.3 Circuit Complexity
We define the circuit complexity classes $\mathsf{AC}^{d}$ and $\mathsf{TC}^{d}$ in the standard way (arora-2009-computational; strobl-etal-2024-formal). A circuit family is a mapping $\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ from input lengths $n$ to circuits that take $n$ inputs. $\mathsf{AC}^{d}$ is the class of language that can be recognized by polynomial-size, fixed-depth circuit families with unbounded-arity AND/OR gates. $\mathsf{TC}^{d}$ is the same class but augmented with unbounded-arity MAJ gates that take a majority vote over their input bits.
It will often be useful to talk about uniform variants of these classes. A $\mathsf{X}$-uniform circuit family obeys that constraint that circuit $C_{n}$ for input size $n$ can be “built” from the input string $1^{n}$ using computation in $\mathsf{X}$. We will consider two standard notions of uniformity: $\mathsf{FO}$ uniformity (which is equivalent to $\mathsf{DLOGTIME}$ uniformity; barrington-1990-uniformity) and the weaker $\mathsf{L}$ uniformity (i.e., log-space uniformity). For a circuit class $\mathsf{XC}$, we will write $\mathsf{A}$-uniform $\mathsf{XC}$ to denote that class constrained to circuit families satisfying $\mathsf{A}$ uniformity. See Appendix˜B for formal definitions of each type of uniformity and strobl-etal-2024-formal for further context.
Finally, we will also use the notion of completeness for circuit classes. Informally, an $\mathsf{X}$-complete problem is a problem in $\mathsf{X}$ to which any other problem in $\mathsf{X}$ can be mapped via some simple reduction $\mathsf{R}$. Whether a problem is complete depends on the notion of reduction used. See Definition˜9 for a more formal definition of reductions, which makes completeness fully defined.